OpenStudy (anonymous):

ratio of 4.35 to 5.8

OpenStudy (jdoe0001):

4.35 <--- notice 2 decimals, thus we use 2 "zeros" for our denominator 5.8 <-- only 1 decimal, thus we use 1 "zero" for our denominator thus \(\bf 4.35\implies \cfrac{435}{100}\qquad \qquad 5.8\implies \cfrac{58}{10}\qquad thus\\ \quad \\ \cfrac{4.35}{5.8}\implies \cfrac{\frac{435}{100}}{\frac{58}{10}}\\ \quad \\ \textit{recall that}\qquad \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\times \cfrac{d}{c}\qquad thus\\ \quad \\ \cfrac{\frac{435}{100}}{\frac{58}{10}}\implies \cfrac{435}{100}\times \cfrac{10}{58}\implies \cfrac{435\times 10}{100\times 58}\)

OpenStudy (jdoe0001):

multiply and simplify the fraction